OpenStudy (anonymous):

Given the function g(x) = 8x - 2, compare and contrast g(-2) and g(4). Choose the statement that is true concerning these two values. The value of g(-2) is larger than the value of g(4). The value of g(-2) is the same as the value of g(4). The value of g(-2) is smaller than the value of g(4). The values of g(-2) and g(4) cannot be compared.

OpenStudy (anonymous):

This is easy, look: It's just telling you to plug in x = -2 and x = 4 into your function: g(x) = 8x - 2 x = -2 g(x) = 8x - 2 g(x) = 8(-2) - 2 g(x) = -16 - 2 g(x) = -18 x = 4 g(x) = 8x - 2 g(x) = 8(4) - 2 g(x) = 32 - 2 g(x) = 30 Can you tell which is the correct answer?

OpenStudy (anonymous):

g(-2) = -18 g(4) = 30

OpenStudy (anonymous):

The answer is C. g(-2) < g(4) Because -18 is less than 30.

OpenStudy (anonymous):

ok thanks :D could you help me out with one more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

The following function defines a recursive sequence. f(0) = -2 f(1) = 8 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -2, 8, -26, -80, … -2, 8, -26, 80, … -2, 12, -44, 180, … -2, -12, -44, -180, … Im not really good at recursive formula, I think its A or C but im not completely sure

OpenStudy (anonymous):

I'm not that good at it either, let me see what I can do.

OpenStudy (anonymous):

Well, look it can't be C or D, because it tells you in the question that the first two terms are -2 and 8.

OpenStudy (anonymous):

ok, so would it be B? Because it goes Negative, Positive, Negative, Postitve?

OpenStudy (anonymous):

Yes, I think so.

OpenStudy (anonymous):

I only have 3 more questions could you help me? Then that would be it

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

The graph below shows a line segment XY. graph of line XY going through ordered pairs negative1 comma 3 and 1 comma negative 1 What is the slope of the line segment XY? -2 -1 over 2 1 over 2 2

OpenStudy (anonymous):

Hold on a minute, let me try this on a graph..

OpenStudy (anonymous):

The answer is -2.

OpenStudy (anonymous):

ok thanks, The first four terms of a sequence are shown below. 6, 2, -2, -6 Which of the following functions best defines this sequence? f(1) = 6, f(n + 1) = f(n) + 4; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) - 2; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) + 2; for n ≥ 1 for this one I believe it is C because in each sequence it is subtracted by 4

OpenStudy (anonymous):

Yeah, I think it's C, too.

OpenStudy (anonymous):

thanks your the best!!!

OpenStudy (anonymous):

have a good day!

OpenStudy (anonymous):

Np.